# Philip ${}^{*\,\sharp}$ Yandex, MIPT &Aleksandr Beznosikov

### v1

@@ Appendix B Memory estimation
In this section, we will examine memory requirements for different projection types using the LLaMA-like architecture as an example and show that RandK, column-wise, and blockwise projections result in approximately the same amount of additional memory for a given density value $\rho$ Appendix A. In contrast, the semi-orthogonal projection matrix (GaLore-like) requires a slightly larger value in this setup. Recall that we follow the setup from Zhao et al. (2024a), where Embeddings, RMSNorms, and Logits remain in the state-full subspace throughout the training, so the projection does not interact with them, and they give the same memory overhead for all projection methods.
@@
Appendix C Optimizer state management
@@
Appendix D LLaMA 7B pre-training results.
In this section, we present the results of pre-training a LLaMA 7b model on the C4 dataset for 120k iterations on 12B tokens. See results in Table 14. We conducted the training in pure bfloat16 with the density $\rho=0.0$. We used learning rate $0.0005$ for state-full optimizer and $0.00015$ for state-free optimizer. However, unlike Zhao et al. (2024a), we didn’t use Adam8bit for state-full parameters but rather Adam, so it may not be an entirely fair comparison. Nevertheless, the results show that FRUGAL has the potential for scaling up to 7B parameter models.
Table 14: Pre-training LLaMA 7B on C4 dataset for 120K steps. Validation perplexity and memory estimate are reported. We also indicate the additional memory overhead introduced by the optimization algorithm.
40K 80K 120K
8-bit Adam 18.09 15.47 14.83
8-bit GaLore 17.94 15.39 14.95
FRUGAL, $\rho=0.0$ 17.56 14.50 13.49
Tokens (B) 5.2 10.5 15.7
@@ Appendix E Convergence Theory
Below, we provide an equivalent formulation of Algorithm 2 that enables us to use the proof of the similar structure to SGDM momentum analyis of Liu et al. (2020).
Algorithm 3 FRUGAL(SGDM, SGD): Equivalent to Algorithm 2 for constant step isze
Input: momentum weight $\beta\in[0,1)$, initialization $x^{1}\in\operatorname*{\mathbb{R}^{d}}$ and $m^{0}=0$, step sizes $\{\alpha_{k}:=\alpha>0\}_{k=1}^{K}$, momentum set $J_{k}\subset[d]$ for $k=1,2\ldots$.
@@
2:     Compute stochastic gradient $\tilde{g}^{k}\leftarrow\nabla f_{\zeta^{k}}(x^{k})$;
3:     Update momentum vector $\tilde{m}_{j}^{k}\leftarrow(1-\beta)\tilde{g}_{j}^{k}+\beta\begin{cases}\tilde{m}_{j}^{k-1}&\text{if }j\in J_{k},\\ 0&\text{otherwise;}\end{cases}$
4:     Update iterate $x^{k+\nicefrac{{1}}{{2}}}\leftarrow x^{k}-\alpha\tilde{m}^{k}$;
5:     $x_{j}^{k+1}\leftarrow\begin{cases}\frac{x_{j}^{k+\nicefrac{{1}}{{2}}}}{1-\beta}-\frac{\beta x^{k}_{j}}{1-\beta}&\text{if }j\notin J_{k+1},\\ x_{j}^{k+\nicefrac{{1}}{{2}}}&\text{otherwise;}\end{cases}$
@@
Lemma 1.
Under Assumption 1, the update vector $\tilde{m}^{k}$ in Algorithm 3 satisfies
@@
Moreover, since $\zeta^{1},\zeta^{2},...,\zeta^{k}$ are independent random variables (item 3 of Assumption 1), we can use conditional expectation to show that $\mathbb{E}\left[(\tilde{g}_{j}^{i_{1}}-g_{j}^{i_{1}})(\tilde{g}_{j}^{i_{2}}-g_{j}^{i_{2}})\right]=0$ for $i_{1}\neq i_{2}$. Therefore,
@@
Lemma 2.
Under Assumption 1, the update vector $\tilde{m}^{k}$ in Algorithm 3 further satisfies
@@
By applying triangle inequality and the smoothness of $f$ (item 1 in Assumption 1), we further have
@@
From Lemma 2, we know that the distance of the non-stochastic momentum from $g^{k}$ is bounded by the weighted sum of past successive iterate differences. Furthermore, the coefficients $a_{k,i}$ decays exponentially in $\beta$.
@@
Proposition 1.
Take Assumption 1. Then, for $z^{k}$ defined in equation 10, we have
@@
where we have applied Lemma 3 in the second step.
@@
Let $k_{j}=k-t^{k-1}_{j}+1$. Then, from Lemma 1 we know that
@@ E.1 Convergence of Algorithm 3
Proposition 2.
Let Assumption 1 hold and $\alpha\leq\frac{1-\beta}{2\sqrt{2}L\sqrt{p^{k}_{\max}}\sqrt{\beta+\beta^{2}}}$ in Algorithm 3. Let $\{c_{i}\}_{i=1}^{\infty}$ in equation 9 be defined by
@@
First of all, by Lemma 2 we know that
@@
Theorem 2.
Let Assumption 1 hold and $\alpha^{k}=\alpha\leq\frac{1-\beta}{L(4-\beta+\beta^{2})}$. Then, the iterates of Algorithm 3 satisfy

### v2

@@ Appendix C Memory estimation
In this section, we will examine memory requirements for different projection types using the LLaMA-like architecture as an example and show that RandK, column-wise, and blockwise projections result in approximately the same amount of additional memory for a given density value $\rho$ Appendix A. In contrast, the semi-orthogonal projection matrix (GaLore-like) requires a slightly larger value in this setup. Recall that we follow the setup from Zhao et al. (2024a), where Embeddings, RMSNorms, and Output layer remain in the state-full subspace throughout the training, so the projection does not interact with them, and they give the same memory overhead for all projection methods.
@@
Appendix D Optimizer state management
@@ Appendix E Convergence Theory
Below, we provide an equivalent formulation of Algorithm 2 that enables us to use the proof of the similar structure to SGDM momentum analyis of (Liu et al., 2020).
Algorithm 3 FRUGAL(SGDM, SGD): Equivalent to Algorithm 2 for constant step size
Input: momentum weight $\beta\in[0,1)$, initialization $x^{1}\in\operatorname*{\mathbb{R}^{d}}$ and $m^{0}=0$, step sizes $\{\alpha_{k}:=\alpha>0\}_{k=1}^{K}$, momentum set $J_{k}\subset[d]$ for $k=1,2\ldots$
@@
2:     Compute stochastic gradient $\tilde{g}^{k}\leftarrow\nabla f_{\zeta^{k}}(x^{k})$
3:     Update momentum vector $\tilde{m}_{j}^{k}\leftarrow(1-\beta)\tilde{g}_{j}^{k}+\beta\begin{cases}\tilde{m}_{j}^{k-1}&\text{if }j\in J_{k},\\ 0&\text{otherwise}\end{cases}$
4:     Update iterate $x^{k+\nicefrac{{1}}{{2}}}\leftarrow x^{k}-\alpha\tilde{m}^{k}$
5:     $x_{j}^{k+1}\leftarrow\begin{cases}\frac{x_{j}^{k+\nicefrac{{1}}{{2}}}}{1-\beta}-\frac{\beta x^{k}_{j}}{1-\beta}&\text{if }j\notin J_{k+1},\\ x_{j}^{k+\nicefrac{{1}}{{2}}}&\text{otherwise}\end{cases}$
@@
Lemma E.2.
Under Assumption 5.1, the update vector $\tilde{m}^{k}$ in Algorithm 3 satisfies
@@
Moreover, since $\zeta^{1},\zeta^{2},...,\zeta^{k}$ are independent random variables (item 3 of Assumption 5.1), we can use conditional expectation to show that $\mathbb{E}\left[(\tilde{g}_{j}^{i_{1}}-g_{j}^{i_{1}})(\tilde{g}_{j}^{i_{2}}-g_{j}^{i_{2}})\right]=0$ for $i_{1}\neq i_{2}$. Therefore,
@@
Lemma E.3.
Under Assumption 5.1, the update vector $\tilde{m}^{k}$ in Algorithm 3 further satisfies
@@
By applying triangle inequality and the smoothness of $f$ (item 1 in Assumption 5.1), we further have
@@
From Lemma E.3, we know that the distance of the non-stochastic momentum from $g^{k}$ is bounded by the weighted sum of past successive iterate differences. Furthermore, the coefficients $a_{k,i}$ decays exponentially in $\beta$.
@@
Proposition E.5.
Take Assumption 5.1. Then, for $z^{k}$ defined in equation 10, we have
@@
where we have applied Lemma E.4 in the second step.
@@
Let $k_{j}=k-t^{k-1}_{j}+1$. Then, from Lemma E.2 we know that
@@ E.1 Convergence of Algorithm 3
Proposition E.6.
Let Assumption 5.1 hold and $\alpha\leq\frac{1-\beta}{2\sqrt{2}L\sqrt{p^{k}_{\max}}\sqrt{\beta+\beta^{2}}}$ in Algorithm 3. Let $\{c_{i}\}_{i=1}^{\infty}$ in equation 9 be defined by
@@
First of all, by Lemma E.3 we know that
@@
Theorem E.7.
Let Assumption 5.1 hold and $\alpha^{k}=\alpha\leq\frac{1-\beta}{L(4-\beta+\beta^{2})}$. Then, the iterates of Algorithm 3 satisfy
@@
Appendix F Limitations
We would also like to acknowledge the limitations of this work. Due to computational constraints, we were unable to conduct experiments on pre-training 7B+ LLMs, which is crucial for understanding the potential of our approach when scaling. Furthermore, our experiments are limited to training language models, although memory-efficient optimization could also be beneficial for training diffusion models. Finally, there may be a better method for selecting the next state-full subspace during the training. We leave the exploration of more sophisticated selection strategies for future work.
Appendix G Simplified algorithms pseudocode
In this section we present the simplified pseudocode of FRUGAL. In Algorithm 4 one can find optimizer steps both for FRUGAL with SVD projection (GaLore-like (Zhao et al., 2024a)) and Block projection (BAdam-like (Luo et al., 2024)).
1:  def svd_or_randk_step(self):
2:    for param in self.params:
3:      grad = param.grad
4:      param_state = self.state[param]
5:      # update projector if necessary
6:      if self.step % self.update_gap == 0:
7:        param_state["projector"] = self.update_proj(grad)
8:      projector = param_state["projector"]
9:      # obtain state-full grad and state-free grad
10:      grad_full = projector.proj_down(grad)
11:      grad_free = grad_full - projector.proj_up(grad_full)
12:      # reset state-full optimizer state if necessary
13:      if self.step % self.update_gap == 0:
14:        param_state["exp_avg"] = torch.zeros_like(grad_full)
15:        param_state["exp_avg_sq"] = torch.zeros_like(grad_full)
16:      # state-full subspace update
17:      self.step += 1
18:      update_full = self.state_full_step(grad_full, param_state)
19:      update_full = projector.proj_up(update_full)
20:      # state-free subspace update
21:      update_free = self.state_free_step(grad_free)
22:      # perform resulting update
23:      update = update_full + update_free
24:      param.add_(update)
25:
26:  def block_step(self):
27:    # change state-full and state-free blocks if necessary
28:    if self.step % self.update_gap == 0:
29:      indices_full = self.update_indices(indices_full)
30:      for idx, param in enumerate(self.params):
31:        grad = param.grad
32:        param_state = self.state[param]
33:        if idx in indices_full:
34:          # reset state-full optimizer state
35:          param_state["exp_avg"] = torch.zeros_like(grad)
36:          param_state["exp_avg_sq"] = torch.zeros_like(grad)
37:          param_state["full_subspace"] = True
38:        else:
39:          # free state-full optimizer state to save memory
40:          param_state.clear()
41:          param_state["full_subspace"] = False
42:    # perform updates
43:    for param in self.params:
44:      grad = param.grad
45:      param_state = self.state[param]
46:      # choose the optimizer depending on the block type
47:      if param_state["full_subspace"]:
48:        update = self.state_full_step(grad, param_state)
49:      else:
50:        update = self.state_free_step(grad)
51:      # perform resulting update
52:      param.add_(update)
Algorithm 4 FRUGAL step pseudocode, PyTorch-like
1:  def state_full_adam_step(self, grad, param_state):
2:    exp_avg = param_state["exp_avg"]
3:    exp_avg_sq = param_state["exp_avg_sq"]
4:    step = self.step
5:    beta1, beta2 = self.betas
6:    exp_avg.mul_(beta1).add_(grad, alpha=1.0-beta1)
7:    exp_avg_sq.mul_(beta2).addcmul_(grad, grad, value=1.0-beta2)
8:    denom = exp_avg_sq.sqrt()
9:    step_size = self.lr_full
10:    if self.correct_bias:
11:      bias_correction1 = 1.0 - beta1 ** step
12:      bias_correction2 = 1.0 - beta2 ** step
13:      step_size = self.lr_full / bias_correction1
14:      bias_correction2_sqrt = math.sqrt(bias_correction2)
15:      denom.div_(bias_correction2_sqrt)
16:    denom.add_(self.eps)
17:    update_full = exp_avg / denom * (-step_size)
18:    return update_full
19:
20:  def state_free_signsgd_step(self, grad):
21:    update_free = -self.lr_free * grad.sign()
22:    return update_free
Algorithm 5 Examples of state-full and state-free steps for Algorithm 4